### v1

@@ 1 Introduction
Graphical models are powerful probabilistic modeling tools for capturing complex relationships among large collections of random variables and have found ample applications in computer vision, natural language processing, signal processing, bioinformatics, and statistics [42]. In this framework, dependencies among random variables are represented by a graph. If this graph is a directed acyclic graph, the graphical model is often referred to as a Baysian Network, while if the graph is undirected, it is often referred to as an undirected graphical model or a Markov Random Field. In this paper, we focus on undirected graphical models (UGMs).
@@
Let $G=(V,E)$ be an undirected graph, where $V,E$ denote node set and edge set of $G$, respectively. In order to define a graphical model, we associate with each node $v\in V$ a random variable $X_{v}$ taking values in some state space ${\mathcal{X}}_{v}$. The edge set $E$ represents dependencies between random variables; that is, for any three distinct nodes $u,v,w\in V$, $X_{u}$ is independent of $X_{v}$ given $X_{w}$, if every path from $u$ to $v$ in $G$ passes through $w$. The notation ${\mathbb{P}}(X_{v}=x_{v})$ corresponds to the probability of the event that the random variable $X_{v}$ takes the value $x_{v}\in{\mathcal{X}}_{v}$. Denote by ${\mathcal{C}}$ the set of maximal cliques in $G$; i.e., the set of cliques that are not properly contained in any other clique of $G$. Recall that a clique $C$ is a subset of $V$ such that $\{u,v\}\in E$ for all $u\neq v\in C$. For each $C\in{\mathcal{C}}$, let us define a nonnegative potential function $\phi_{C}(x_{C})$, where $x_{C}$ is the vector consisting of $x_{v}$, $v\in C$. In this paper, we assume ${\mathcal{X}}_{v}\in\{0,1\}$ for all $v\in V$, henceforth referred to as a binary UGM. It can be shown that the joint probability mass function for a binary UGM is given by:
@@
where $Z$ is a normalization constant given by $Z:=\sum_{x}{\prod_{C\in{\mathcal{C}}}\phi_{C}(x_{C})}$. The order of a UGM is defined as the size of a largest clique $C\in{\mathcal{C}}$ minus one. Due to their simplicity, first-order UGMs, also known as pair-wise models, are the most popular UGMs. However, to model more complex interactions among random variables it is essential to study higher-order UGMs.
@@
Given some noisy observation $Y_{v}$, $v\in V$, we would like to recover the ground truth $X_{v}$, $v\in V$, whose probability mass function is described by a binary UGM defined by (1). By definition, the maximum a posteriori (MAP) estimator maximizes the probability of recovering the ground truth. In the following, we denote by ${\mathbb{P}}[X|Y]$ the probability that $X_{v}$, $v\in V$ is the ground truth given that $Y_{v}$, $v\in V$ is observed. Hence, we are interested in solving the following optimization problem:
@@
By Bayes’ theorem and monotonicity of the log function, we deduce that
@@
${\mathbb{P}}[Y\mid X]=\prod_{v\in V}{\mathbb{P}}\left[y_{v}\mid x_{v}\right]=\prod_{v\in V:y_{v}=0}{p^{x_{v}}(1-p)^{1-x_{v}}}\prod_{v\in V:y_{v}=1}{p^{1-x_{v}}(1-p)^{x_{v}}}.$
@@
where $\alpha$ is a positive parameter that along with the remaining parameters $c_{e}$, $e\in P(C)$, $C\in{\mathcal{C}}$ should be learned from the data.
@@
The literature on inference in UGMs is mostly focused on first-order UGMs; i.e., the case where $|C|=2$ for all $C\in{\mathcal{C}}$. For first-order binary UGMs, Problem 6 simplifies to an unconstrained binary quadratic optimization problem, which is NP-hard in general. The most popular methods to tackle this inference problem are belief propagation [41, 21], which is a message passing algorithm, and graph cut algorithms [32, 4, 5, 31]. Utilizing higher-order UGMs is essential for capturing more complex interactions among random variables. Yet, their study has been fairly limited due to the complexity of solving Problem 6 in its full generality. In fact, almost all existing studies considering higher-order UGMs tackle the inference problem by first reducing it to a binary quadratic optimization problem through the introduction of auxiliary variables and subsequently employing graph cut algorithms to solve the quadratic optimization problem [22, 38, 26, 27]. In [20], the authors consider the inference problem for a higher-order binary UGM arising from error-correcting decoding problem and propose a linear programming (LP) relaxation for this problem. In [19], the authors analyze the performance of the LP relaxation of [20] theoretically, hence establishing the effectiveness of the LP relaxation for decoding low-density-parity-check codes. In [8] the authors consider a third-order binary UGM for a simplified image restoration problem and propose an LP relaxation for this problem.
@@
In spite of its ample applications, existing results for inference in higher-order binary UGMs are rather scarce. In this paper, by building upon recent theoretical and algorithmic developments for binary polynomial optimization [10, 12, 11, 13, 16, 14], we present strong LP relaxations for Problem 6 in its full generality. We prove that the proposed LPs are stronger than the existing LPs for this class of problems and can be solved efficiently using off-the-shelf LP solvers. We consider two important applications of inference in higher-order binary UGMs; namely image restoration, a popular application in computer vision, and decoding error-correcting codes, a central problem in information theory. Via an extensive computational study, we show that a simple LP relaxation that we refer to as the “clique LP” is often sharp for image restoration problems. The decoding problem on the other hand turns out to be a difficult problem and while the proposed clique LP outperforms the only existing LP relaxation for this problem [19], the improvement is rather small.
@@ 2 Linear programming relaxations
With the objective of constructing LP relaxations for Problem 6, following a common practice in nonconvex optimization, we start by linearizing its objective function. For notational simplicity, henceforth we denote variables $x_{v}$, $v\in V$ by $z_{v}$, $v\in V$. Define $\bar{P}(C):=P(C)\setminus(C\cup\{\emptyset\})$ for all $C\in{\mathcal{C}}$. Define an auxiliary variable $z_{e}:=\prod_{v\in e}{z_{v}}$ for all $e\in\bar{P}(C)$ and for all $C\in{\mathcal{C}}$. Then an equivalent reformulation of Problem 6 in a lifted space of variables is given by:
@@
We refer to the convex hull of $S({\mathcal{G}})$ as the multilinear polytope and denote it by $\text{MP}({\mathcal{G}})$. Henceforth, we refer to the hypergraph ${\mathcal{G}}$ associated with Problem B-UGM as the UGM hypergraph. To obtain an LP relaxation for Problem B-UGM, it suffices to construct a polyhedral relaxation for the multilinear set $S({\mathcal{G}})$. Notice that the rank of a UGM hypergraph equals the size of the largest clique $C$ in the corresponding UGM; this number in turn is always quite small in practice. This key property enables us to obtain strong and yet cheap relaxations for the multilinear polytope of a UGM hypergraph. In the remainder of this section, we briefly review existing LP relaxations for Problem B-UGM; subsequently, we propose new LP relaxations for it.
@@ 2.1 The standard linearization
In [11, 6], the authors prove that $\text{MP}^{\text{LP}}({\mathcal{G}})=\text{MP}({\mathcal{G}})$ if and only if ${\mathcal{G}}$ is a Berge-acyclic hypergraph; i.e., the most restrictive type of hypergraph acyclicity [18]. However, a UGM hypergraph is not Berge-acyclic and indeed our numerical experiments indicate that Problem stdLP leads to very weak upper bounds for Problem B-UGM.
@@ 2.2 The flower relaxation
If $|T|=1$, flower inequalities simplify to two-link inequalities introduced in [8]. We define the flower relaxation $\text{MP}^{F}({\mathcal{G}})$ as the polytope obtained by adding all flower inequalities centered at each edge of ${\mathcal{G}}$ to $\text{MP}^{\text{LP}}({\mathcal{G}})$. In [16], the authors prove that while the separation problem over the flower relaxation is ${\mathcal{N}P}$-hard for general hypergraphs, it can be solved in polynomial time for hypergraphs with fixed rank. As we discussed before, in case of a UGM hypergraph, it is reasonable to assume that $r$ is fixed. In fact, as we show next, for a UGM hypergraph, it suffices to include only a small number of flower inequalities in the flower relaxation.
@@
$\gamma(e_{0},e_{k},k\in T):=\bigcup_{k\in T}{e_{k}}\cup e_{0}\subseteq C,\quad{\rm for\;some}\;C\in{\mathcal{C}},$ (11)
@@
Let $e_{0},e_{k}$, $k\in T$ be edges of ${\mathcal{G}}$ satisfying condition (9) but not satisfying condition (11). Then the following flower inequality is present in $\text{MP}^{F}({\mathcal{G}})$:
@@
First consider the case where $e_{j}\supset e_{0}$ for some $j\in T$; then by (9) we must have $|T|=1$, implying that $\gamma(e_{0},e_{j})=e_{j}$ and hence condition (11) holds. Henceforth, suppose that $e_{k}\not\supset e_{0}$ for all $k\in T$. Notice that if $e_{k}\subset e_{0}$ for all $k\in T$, then we have $\gamma(e_{0},e_{k},k\in T)=e_{0}$ and condition (11) is trivially satisfied. Denote by $T^{\prime}$ the nonempty set containing all $k\in T$ satisfying $e_{k}\setminus e_{0}\neq\emptyset$. Define $\tilde{e}_{k}=e_{k}\cap e_{0}$ for all $k\in T^{\prime}$. By definition of a UGM hypergraph and condition (9), we have $\tilde{e}_{k}\in{\mathcal{E}}$ for all $k\in T^{\prime}$. Hence the following flower inequalities are also present in $\text{MP}^{F}({\mathcal{G}})$:
@@
where we used $e_{0}\setminus\cup_{k\in T}{e_{k}}=e_{0}\setminus(\cup_{k\in T\setminus T^{\prime}}{e_{k}}\cup\cup_{k\in T^{\prime}}{\tilde{e}_{k}})$. First, it is simple to verify that condition (11) is satisfied for inequalities (13) and (14). Second, summing up inequalities (13) and (14), we obtain inequality (12), implying its redundancy, and this completes the proof. ∎
@@ 2.3 The running intersection relaxation
where we define $z_{\emptyset}=0$, and
$\omega=\Big{|}e_{0}\setminus\bigcup_{k\in T}e_{k}\Big{|}+\Big{|}\Big{\{}k\in T:N(e_{0}\cap e_{k})=\emptyset\Big{\}}\Big{|}.$
Notice that by letting $w_{k}=\emptyset$ for all $k\in T$, the running intersection inequality (17) simplifies to the flower inequality (10). Consider $e_{0},e_{k}$, $k\in T$ such that $\tilde{E}$ has the running intersection property and $N(e_{0}\cap e_{k})\neq\emptyset$ for some $k\in T$. Then any running intersection inequality centered at $e_{0}$ with neighbours $e_{k}$, $k\in T$ satisfying $w_{k}\neq\emptyset$ for some $k\in T$ together with $z_{w_{k}}\leq 1$ imply the flower inequality centered at $e_{0}$ with neighbours $e_{k}$, $k\in T$. However, in general flower inequalities are not implied by running intersection inequalities, since for flower inequalities we do not require the set $\tilde{E}$ to have the running intersection property.
@@
In [16], the authors prove that for hypergraphs with fixed rank, the separation problem over running intersection inequalities can be solve in polynomial time. Our computational results indicate that the running LP is significantly stronger than the flower LP. However, the added strength often comes at the cost of a rather significant increase in CPU time.
@@ 2.4 The clique relaxation
Proposition 1.
(Theorem 2 in [39]) Let ${\mathcal{G}}_{C}$ be a complete hypergraph with node set $C$. Then the multilinear polytope $\text{MP}({\mathcal{G}}_{C})$ is given by
@@
In Sections 3 and 4 we show that the clique LP returns a binary solution in many cases of practical interest. We next present a theoretical justification of this fact. That is, we show that all inequalities defining facets of the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}})$ are facet-defining for the multilinear polytope $\text{MP}({\mathcal{G}})$ as well. To this end, we make use of a zero-lifting operation for the multilinear polytope that was introduced in [10]. Let ${\mathcal{G}}=(V,{\mathcal{E}})$ be a hypergraph. Then the hypergraph ${\mathcal{G}}^{\prime}=(V^{\prime},{\mathcal{E}}^{\prime})$ is a partial hypergraph of ${\mathcal{G}}$ if $V^{\prime}\subseteq V$ and ${\mathcal{E}}^{\prime}\subseteq{\mathcal{E}}$. The following lemma [10], provides a sufficient condition under which a facet-defining inequality for $\text{MP}({\mathcal{G}}^{\prime})$ is also facet-defining for $\text{MP}({\mathcal{G}})$.
Lemma 3.
(Corollary 4 in [10]) Let the complete hypergraph ${\mathcal{G}}^{\prime}=(V^{\prime},{\mathcal{E}}^{\prime})$ be a partial hypergraph of ${\mathcal{G}}=(V,{\mathcal{E}})$. Then all facet-defining inequalities for $\text{MP}({\mathcal{G}}^{\prime})$ are facet-defining for $\text{MP}({\mathcal{G}})$ if and only if there exists no edge $e\in{\mathcal{E}}$ such that $e\supset V^{\prime}$.
@@
Since $C\in{\mathcal{C}}$ is a maximal clique of the UGM, by definition, there exists no edge $e\in{\mathcal{E}}$ that strictly contains $C$. Hence assumptions of Lemma 3 are satisfied and the result follows. ∎
By construction, for a general UGM hypergraph ${\mathcal{G}}$ we have $\text{MP}^{\text{LP}}({\mathcal{G}})\subset\text{MP}^{F}({\mathcal{G}})\subset\text{MP}^{\rm RI}({\mathcal{G}})$. The next result indicates that the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}})$ is the strongest relaxation of $\text{MP}({\mathcal{G}})$ introduced so far.
@@
Let ${\mathcal{G}}=\cup_{C\in{\mathcal{C}}}{{\mathcal{G}}_{C}}$ be a UGM hypergraph of rank $r\geq 3$, where ${\mathcal{C}}$ denotes the set of maximal clique of the binary UGM. Then $\text{MP}^{\rm cl}({\mathcal{G}})\subsetneq\text{MP}^{\rm RI}({\mathcal{G}})$.
@@
$\sum_{\begin{subarray}{c}W\subseteq C:\\ |W|\;{\rm even}\end{subarray}}{z_{W}}-\sum_{\begin{subarray}{c}W\subseteq C:\\ |W|\;{\rm odd}\end{subarray}}{z_{W}}\geq 0,$
which is clearly not present in $\text{MP}^{\rm RI}({\mathcal{G}})$, if $|C|\geq 3$. ∎
Next we provide a necessary and sufficient condition under which the clique relaxation coincides with the multilinear polytope. To this end we make use of two tools which we present first. First, we outline a sufficient condition for decomposability of multilinear sets [12]. Let ${\mathcal{G}}=(V,{\mathcal{E}})$ be a hypergraph and let ${\mathcal{G}}^{\prime}=(V^{\prime},{\mathcal{E}}^{\prime})$ be a partial hypergraph of ${\mathcal{G}}$. Then ${\mathcal{G}}^{\prime}$ is a section hypergraph of ${\mathcal{G}}$ induced by $V^{\prime}$, if =${\mathcal{E}}^{\prime}=\{e\in{\mathcal{E}}:e\subseteq V^{\prime}\}$. In the following, given hypergraphs ${\mathcal{G}}_{1}=(V_{1},{\mathcal{E}}_{1})$ and ${\mathcal{G}}_{2}=(V_{2},{\mathcal{E}}_{2})$, we denote by ${\mathcal{G}}_{1}\cap{\mathcal{G}}_{2}$ the hypergraph $(V_{1}\cap V_{2},{\mathcal{E}}_{1}\cap{\mathcal{E}}_{2})$, and we denote by ${\mathcal{G}}_{1}\cup{\mathcal{G}}_{2}$, the hypergraph $(V_{1}\cup V_{2},{\mathcal{E}}_{1}\cup{\mathcal{E}}_{2})$. Finally, we say that the multilinear set $S({\mathcal{G}})$ is decomposable into $S({\mathcal{G}}_{1})$ and $S({\mathcal{G}}_{2})$ if the system comprised of the description of $\text{MP}({\mathcal{G}}_{1})$ and the description of $\text{MP}({\mathcal{G}}_{2})$, is the description of $\text{MP}({\mathcal{G}})$.

### v2

@@ 1 Introduction
Graphical models are powerful probabilistic modeling tools for capturing complex relationships among large collections of random variables and have found ample applications in computer vision, natural language processing, signal processing, bioinformatics, and statistics [44]. In this framework, dependencies among random variables are represented by a graph. If this graph is a directed acyclic graph, the graphical model is often referred to as a Bayesian Network, while if the graph is undirected, it is often referred to as an Undirected Graphical Model (UGM) or a Markov Random Field. In this paper, we focus on UGMs.
@@
Let $G=(V,E)$ be an undirected graph, where $V,E$ denote node set and edge set of $G$, respectively. In order to define a graphical model, we associate with each node $v\in V$ a random variable $X_{v}$ taking values in some state space ${\mathcal{X}}_{v}$. The edge set $E$ represents dependencies between random variables; that is, for any three distinct nodes $u,v,w\in V$, $X_{u}$ is independent of $X_{v}$ given $X_{w}$, if every path from $u$ to $v$ in $G$ passes through $w$. The notation ${\mathbb{P}}(X_{v}=x_{v})$ corresponds to the probability of the event that the random variable $X_{v}$ takes the value $x_{v}\in{\mathcal{X}}_{v}$. Denote by ${\mathcal{C}}$ the set of maximal cliques in $G$; i.e., the set of cliques that are not properly contained in any other clique of $G$. Recall that a clique $C$ is a subset of $V$ such that $\{u,v\}\in E$ for all $u\neq v\in C$. For each $C\in{\mathcal{C}}$, let us define a nonnegative potential function $\phi_{C}(x_{C})$, where $x_{C}$ is the vector consisting of $x_{v}$, $v\in C$. In this paper, we assume ${\mathcal{X}}_{v}=\{0,1\}$ for all $v\in V$, henceforth referred to as a binary UGM. It can be shown that the joint probability mass function for a binary UGM is given by:
@@
where $Z$ is a normalization constant given by $Z:=\sum_{x}{\prod_{C\in{\mathcal{C}}}\phi_{C}(x_{C})}$. The order of a UGM is defined as the size of a largest clique $C\in{\mathcal{C}}$ minus one. Due to their simplicity, first-order UGMs, also known as pair-wise models, are the most popular UGMs. However, to model more complex interactions among random variables, it is essential to study higher-order UGMs.
@@
Given some noisy observation $Y_{v}$, $v\in V$, we would like to recover the ground truth $X_{v}$, $v\in V$, whose probability mass function is described by a binary UGM defined by (1). By definition, the maximum a posteriori (MAP) estimator maximizes the probability of recovering the ground truth. In the following, we denote by ${\mathbb{P}}[X|Y]$ the probability that $X_{v}$, $v\in V$ is the ground truth, given that $Y_{v}$, $v\in V$ is observed. Hence, we are interested in solving the following optimization problem:
@@
Notice that ${\mathbb{P}}[Y]>0$, and does not depend on $x$. Hence, by Bayes’ theorem and monotonicity of the log function, we deduce that
@@
${\mathbb{P}}[Y\mid X]=\prod_{v\in V}{\mathbb{P}}\left[y_{v}\mid x_{v}\right]=\prod_{v\in V:y_{v}=0}{p^{x_{v}}(1-p)^{1-x_{v}}}\prod_{v\in V:y_{v}=1}{p^{1-x_{v}}(1-p)^{x_{v}}},$
which in turn implies that
$\log({\mathbb{P}}[Y\mid X])$ $=\log\Big{(}\frac{1-p}{p}\Big{)}\Big{(}\sum_{v\in V:y_{v}=1}{x_{v}}-\sum_{v\in V:y_{v}=0}{x_{v}}\Big{)}+\Big{|}\{v\in V:y_{v}=1\}\Big{|}\log p$
+|$+\Big{|}\{v\in V:y_{v}=0\}\Big{|}\log(1-p).$
@@
where $\alpha$ is a positive parameter that along with the remaining parameters $c_{e}$, $e\in P(C)$, $C\in{\mathcal{C}}$ are learned from the data.
@@
The literature on inference in UGMs is mostly focused on first-order UGMs; i.e., the case where $|C|=2$ for all $C\in{\mathcal{C}}$. For first-order binary UGMs, Problem 6 simplifies to an unconstrained binary quadratic optimization problem, which is NP-hard in general. The most popular methods to tackle this inference problem are belief propagation [43, 21], which is a message passing algorithm, and graph cut algorithms [33, 4, 5, 32]. Moreover, constant-factor approximation algorithms are available for this problem class [26, 37]. Utilizing higher-order UGMs is essential for capturing more complex interactions among random variables. Yet, their study has been fairly limited due to the complexity of solving Problem 6 in its full generality. In fact, almost all existing studies considering higher-order UGMs tackle the inference problem by first reducing it to a binary quadratic optimization problem through the introduction of auxiliary variables and subsequently employing graph cut algorithms to solve the quadratic optimization problem [22, 40, 27, 28]. In [20], the authors consider the inference problem for a higher-order binary UGM arising from the error-correcting decoding problem and propose a linear programming (LP) relaxation for this problem. In [19], the authors analyze the performance of the LP relaxation of [20] theoretically, hence establishing the effectiveness of the LP relaxation for decoding low-density-parity-check codes. In [8] the authors consider a third-order binary UGM for a simplified image restoration problem and propose an LP relaxation for this problem.
@@
In spite of its ample applications, the existing results for inference in higher-order binary UGMs are rather scarce. In this paper, by building upon recent theoretical and algorithmic developments for binary polynomial optimization [10, 12, 11, 13, 16, 14], we present strong LP relaxations for Problem 6 in its full generality. We prove that the proposed LPs are stronger than the existing LPs for this class of problems and can be solved efficiently using off-the-shelf LP solvers. We consider two important applications of inference in higher-order binary UGMs; namely image restoration, a popular application in computer vision, and decoding error-correcting codes, a central problem in information theory. Via an extensive computational study, we show that a simple LP relaxation that we refer to as the “clique LP” is often sharp for image restoration problems. The decoding problem on the other hand turns out to be a difficult problem and while the proposed clique LP outperforms the only existing LP relaxation for this problem [19], the improvement is rather small.
@@ 2 Linear programming relaxations
With the objective of constructing LP relaxations for Problem 6, following a common practice in nonconvex optimization, we start by linearizing its objective function. Define $\bar{P}(C):=P(C)\setminus(C\cup\{\emptyset\})$ for all $C\in{\mathcal{C}}$. For notational simplicity, henceforth we denote variables $x_{v}$, $v\in V$ by $z_{v}$, $v\in V$. Define an auxiliary variable $z_{e}:=\prod_{v\in e}{z_{v}}$ for all $e\in\bar{P}(C)$ and for all $C\in{\mathcal{C}}$. Then an equivalent reformulation of Problem 6 in a lifted space of variables is given by:
@@
We refer to the convex hull of $S({\mathcal{G}})$ as the multilinear polytope and denote it by $\text{MP}({\mathcal{G}})$. Henceforth, we refer to the hypergraph ${\mathcal{G}}$ associated with Problem B-UGM as an UGM hypergraph. To obtain an LP relaxation for Problem B-UGM, it suffices to construct a polyhedral relaxation for the multilinear set $S({\mathcal{G}})$. Notice that the rank of a UGM hypergraph equals the size of the largest clique $C$ in the corresponding UGM; this number in turn is always quite small in practice. This key property enables us to obtain strong and yet cheaply computable relaxations for the multilinear polytope of a UGM hypergraph. In the remainder of this section, we briefly review existing LP relaxations for Problem B-UGM; subsequently, we propose new LP relaxations for it.
@@ 2.1 The standard linearization
In [11, 6], the authors prove that $\text{MP}^{\text{LP}}({\mathcal{G}})=\text{MP}({\mathcal{G}})$ if and only if ${\mathcal{G}}$ is a Berge-acyclic hypergraph; i.e., the most restrictive type of acyclic hypergraphs [18]. However, a UGM hypergraph is not Berge-acyclic and indeed our numerical experiments indicate that Problem stdLP leads to very weak upper bounds for Problem B-UGM.
@@ 2.2 The flower relaxation
If $|T|=1$, flower inequalities simplify to two-link inequalities introduced in [8]. We define the flower relaxation $\text{MP}^{F}({\mathcal{G}})$ as the polytope obtained by adding all flower inequalities centered at each edge of ${\mathcal{G}}$ to $\text{MP}^{\text{LP}}({\mathcal{G}})$. In [16], the authors prove that while the separation problem over the flower relaxation is ${\mathcal{N}P}$-hard for general hypergraphs, it can be solved in polynomial time for hypergraphs with fixed rank. As we discussed before, in case of a UGM hypergraph, it is reasonable to assume that the rank is fixed. In fact, as we show next, for a UGM hypergraph, it suffices to include only a small number of flower inequalities in the flower relaxation.
@@
$\bigcup_{k\in T}{e_{k}}\cup e_{0}\subseteq C,\quad{\rm for\;some}\;C\in{\mathcal{C}},$ (11)
@@
By construction, $\text{MP}^{F}({\mathcal{G}})\subseteq\text{MP}^{F^{\prime}}({\mathcal{G}})$. Hence, it suffices to show that $\text{MP}^{F^{\prime}}({\mathcal{G}})\subseteq\text{MP}^{F}({\mathcal{G}})$. To this end, we show that any flower inequality not satisfying condition (11) is implied by some flower inequalities that satisfy this condition. Let $e_{0},e_{k}$, $k\in T$ be edges of ${\mathcal{G}}$ satisfying condition (9) but not satisfying condition (11). Then the following flower inequality is present in $\text{MP}^{F}({\mathcal{G}})$:
@@
Define $\gamma(e_{0},e_{k},k\in T):=\bigcup_{k\in T}{e_{k}}\cup e_{0}$. First consider the case where $e_{j}\supset e_{0}$ for some $j\in T$; then by (9) we must have $|T|=1$, implying that $\gamma(e_{0},e_{j})=e_{j}$ and hence condition (11) holds. Henceforth, suppose that $e_{k}\not\supset e_{0}$ for all $k\in T$. Notice that if $e_{k}\subset e_{0}$ for all $k\in T$, then we have $\gamma(e_{0},e_{k},k\in T)=e_{0}$ and condition (11) is trivially satisfied. Denote by $T^{\prime}$ the nonempty set containing all $k\in T$ satisfying $e_{k}\setminus e_{0}\neq\emptyset$. Define $\tilde{e}_{k}=e_{k}\cap e_{0}$ for all $k\in T^{\prime}$. By definition of a UGM hypergraph and condition (9), we have $\tilde{e}_{k}\in{\mathcal{E}}$ for all $k\in T^{\prime}$. Hence the following flower inequalities are also present in $\text{MP}^{F}({\mathcal{G}})$:
@@
where we used the identity $e_{0}\setminus\cup_{k\in T}{e_{k}}=e_{0}\setminus((\cup_{k\in T\setminus T^{\prime}}{e_{k}})\cup(\cup_{k\in T^{\prime}}{\tilde{e}_{k}}))$. First, it is simple to verify that condition (11) is satisfied for inequalities (13) and (14). Second, summing up inequalities (13) and (14), we obtain inequality (12), implying its redundancy, and this completes the proof. ∎
@@ 2.3 The running intersection relaxation
where we define $z_{\emptyset}:=0$, and
$\omega:=\Big{|}e_{0}\setminus\bigcup_{k\in T}e_{k}\Big{|}+\Big{|}\Big{\{}k\in T:N(e_{0}\cap e_{k})=\emptyset\Big{\}}\Big{|}.$
Notice that by letting $w_{k}=\emptyset$ for all $k\in T$, the running intersection inequality (17) simplifies to the flower inequality (10). Consider $e_{0},e_{k}$, $k\in T$ such that $\tilde{E}$ has the running intersection property and $N(e_{0}\cap e_{k})\neq\emptyset$ for some $k\in T$. Then any running intersection inequality centered at $e_{0}$ with neighbors $e_{k}$, $k\in T$ satisfying $w_{k}\neq\emptyset$ for some $k\in T$ together with $z_{w_{k}}\leq 1$ imply the flower inequality centered at $e_{0}$ with neighbors $e_{k}$, $k\in T$. However, in general flower inequalities are not implied by running intersection inequalities, since for flower inequalities we do not require the set $\tilde{E}$ to have the running intersection property.
@@
In [16], the authors prove that for hypergraphs with fixed rank, the separation problem over running intersection inequalities can be solved in polynomial time. Our computational results indicate that the running LP is significantly stronger than the flower LP. However, the added strength often comes at the cost of a rather significant increase in CPU time.
@@ 2.4 The clique relaxation
Proposition 1 (Theorem 2 in [41]).
Let ${\mathcal{G}}_{C}$ be a complete hypergraph with node set $C$. Then the multilinear polytope $\text{MP}({\mathcal{G}}_{C})$ is given by
@@
In Sections 3 and 4 we show that the clique LP returns a binary solution in many cases of practical interest. We next present a theoretical justification of this fact. That is, we show that all inequalities defining facets of the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}})$ are facet-defining for the multilinear polytope $\text{MP}({\mathcal{G}})$ as well. To this end, we make use of a zero-lifting operation for the multilinear polytope that was introduced in [10]. Let ${\mathcal{G}}=(V,{\mathcal{E}})$ be a hypergraph. Then the hypergraph ${\mathcal{G}}^{\prime}=(V^{\prime},{\mathcal{E}}^{\prime})$ is a partial hypergraph of ${\mathcal{G}}$ if $V^{\prime}\subseteq V$ and ${\mathcal{E}}^{\prime}\subseteq{\mathcal{E}}$. The following lemma provides a sufficient condition under which a facet-defining inequality for $\text{MP}({\mathcal{G}}^{\prime})$ is also facet-defining for $\text{MP}({\mathcal{G}})$.
Lemma 3 (Corollary 4 in [10]).
Let the complete hypergraph ${\mathcal{G}}^{\prime}=(V^{\prime},{\mathcal{E}}^{\prime})$ be a partial hypergraph of ${\mathcal{G}}=(V,{\mathcal{E}})$. Then all facet-defining inequalities for $\text{MP}({\mathcal{G}}^{\prime})$ are facet-defining for $\text{MP}({\mathcal{G}})$ if and only if there exists no edge $e\in{\mathcal{E}}$ such that $e\supset V^{\prime}$.
@@
Since $C\in{\mathcal{C}}$ is a maximal clique of the UGM, by definition, there exists no edge $e\in{\mathcal{E}}$ that strictly contains $C$. Hence, the assumptions of Lemma 3 are satisfied and the result follows. ∎
By construction, for a general UGM hypergraph ${\mathcal{G}}$ we have $\text{MP}^{\rm RI}({\mathcal{G}})\subset\text{MP}^{F}({\mathcal{G}})\subset\text{MP}^{\rm\text{LP}}({\mathcal{G}})$. The next result indicates that the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}})$ is the strongest relaxation introduced so far.
@@
Let ${\mathcal{G}}=\cup_{C\in{\mathcal{C}}}{{\mathcal{G}}_{C}}$ be a UGM hypergraph of rank $r\geq 3$, where ${\mathcal{C}}$ denotes the set of maximal clique of the binary UGM. Then $\text{MP}^{\rm cl}({\mathcal{G}})\subset\text{MP}^{\rm RI}({\mathcal{G}})$.
@@
$\sum_{\begin{subarray}{c}W\subseteq C:\\ |W|\;{\rm odd}\end{subarray}}{z_{W}}-\sum_{\begin{subarray}{c}W\subseteq C:\\ |W|\;{\rm even}\end{subarray}}{z_{W}}\leq 0,$
which is not present in $\text{MP}^{\rm RI}({\mathcal{G}})$, if $|C|\geq 3$. To see this, note that in the above inequality, the cardinality of the set of variables with negative coefficients, i.e., $\{W\subseteq C:|W|\;{\rm even},|W|\geq 2\}$ is lower bounded by $\binom{|C|}{2}$ which is larger than one, if $|C|\geq 3$. However, in a running intersection inequality (17), there is only one variable $z_{e}$, for some $e\in{\mathcal{E}}$ with a negative coefficient. ∎
Next we provide a sufficient condition under which the clique relaxation coincides with the multilinear polytope. To this we make use of a sufficient condition for decomposability of multilinear sets [12]. Let ${\mathcal{G}}=(V,{\mathcal{E}})$ be a hypergraph and let ${\mathcal{G}}^{\prime}=(V^{\prime},{\mathcal{E}}^{\prime})$ be a partial hypergraph of ${\mathcal{G}}$. Then ${\mathcal{G}}^{\prime}$ is a section hypergraph of ${\mathcal{G}}$ induced by $V^{\prime}$, if =${\mathcal{E}}^{\prime}=\{e\in{\mathcal{E}}:e\subseteq V^{\prime}\}$. Given hypergraphs ${\mathcal{G}}_{1}=(V_{1},{\mathcal{E}}_{1})$ and ${\mathcal{G}}_{2}=(V_{2},{\mathcal{E}}_{2})$, we denote by ${\mathcal{G}}_{1}\cap{\mathcal{G}}_{2}$ the hypergraph $(V_{1}\cap V_{2},{\mathcal{E}}_{1}\cap{\mathcal{E}}_{2})$, and we denote by ${\mathcal{G}}_{1}\cup{\mathcal{G}}_{2}$, the hypergraph $(V_{1}\cup V_{2},{\mathcal{E}}_{1}\cup{\mathcal{E}}_{2})$. Finally, we say that the multilinear set $S({\mathcal{G}})$ is decomposable into $S({\mathcal{G}}_{1})$ and $S({\mathcal{G}}_{2})$ if the system comprised of the description of $\text{MP}({\mathcal{G}}_{1})$ and the description of $\text{MP}({\mathcal{G}}_{2})$, is the description of $\text{MP}({\mathcal{G}})$.